OpenStudy (anonymous):

can someone help me what is x divided by cuberoot3

OpenStudy (amistre64):

its an expression, or a rule .....

OpenStudy (anonymous):

we have to rationalize the denominator. for some reason the answer i get is wrong idk why

OpenStudy (amistre64):

youre most likely trying to use a square root principle instead of a cube root

OpenStudy (amistre64):

\[\Large\sqrt[3]{3}*\sqrt[3]{3*3}=\sqrt[3]{3*3*3}=\sqrt[3]{3^3}=3\]

OpenStudy (anonymous):

Can u explain where u go the extra 3's from

OpenStudy (amistre64):

to "rationalize" means we want to get rid of a radical ... teh radical you have under there is: cbrt(3) a cbrt and a ^3 cancel each other\[\Large \sqrt[3]{a^3}=a\] as such, you only have a 3^1 in there and need 2 more 3s to undo the radical .... hence, 3^2

OpenStudy (amistre64):

im not sure how familiar you are with changing radicals to exponents, which would be one way to approach this: \[\frac{x}{3^{1/3}}*\frac{3^{2/3}}{3^{2/3}}=\frac{x~3^{2/3}}{3}\]

OpenStudy (anonymous):

oh ok got it thanx!

OpenStudy (amistre64):

youre welcome :)